OpenStudy (anonymous):

Can you determine the factors of x^2 - 8x - 12?

OpenStudy (anonymous):

K so

OpenStudy (anonymous):

any polynomial in form ax^2+bx+c can be factored by find the two numbers that multiply to give you c and add to give you b

OpenStudy (anonymous):

1:SPLIT THE MID FACTOR X^2+8X+12 EQUALS X^2+6X +2X+12 [12=6*2 8=6+2] 2:TAKE COMMON FACTORS FROM 1ST TWO AND LAST TWO FACTORS X^2+6X +2X+12 EQUALS X(X+6)+2(X+6) 3:TAKE (X+6) AS COMMON FACTOR X(X+6)+2(X+6) EQUALS (X+6){X+2} PS:WHILE SOLVING QUADRATIC EQNS IN THIS FORM WHERE X^2 HAS NO COEEFICIENT U CAN DIRECTLY WRITE THE SOLN AS (X+6)(X+2)

OpenStudy (anonymous):

so what two numbers multily together to give you -12 and add to give you -8

OpenStudy (anonymous):

woah marco way to complicated

OpenStudy (anonymous):

You have to use the quadratic equation.

OpenStudy (anonymous):

guys.... you are making this so much more complicated then it needs to be

OpenStudy (anonymous):

k im confused

OpenStudy (anonymous):

Oh yeah that doesnt factor nicely

OpenStudy (anonymous):

k ik its either (x - 6)(x + 2) or (x + 6)(x - 2) but idk which one

OpenStudy (anonymous):

The way I mentioned actually fails in this case

OpenStudy (anonymous):

so make use of the quadratic formula

OpenStudy (anonymous):

but which one is (x - 6)(x + 2) or (x + 6)(x - 2)

OpenStudy (anonymous):

Neither, it must be solved using the quadratic equation.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Use b^2-4ac to determine whether it can be factored nicely if not use the quadratic formula